OpenStudy (anonymous):

a coin is tossed and a standard number cube is rolled. What is the probability that the coin shows heads and the number cube shows an even number?

OpenStudy (anonymous):

What is the probability that the coin will show heads?

OpenStudy (anonymous):

Rena what is the probability of the coin landing on heads?

OpenStudy (anonymous):

I cant help if you dont answer

OpenStudy (anonymous):

\[P(heads)=\frac{1}{2}\]\[P(even)=\frac{3}{6}\]\[P(evenheads)=P(even) \times P(heads)\]\[P(evenheads)=\frac{1}{2} \times \frac{3}{6} = \frac{1}{4}\] or 25%

OpenStudy (anonymous):

Of course dont teach him how to do it just show the work with the answer. It really teaches him how to do math when you do that

OpenStudy (anonymous):

He asked what the probability was, not how to find the probability. I answered the question that was asked.

OpenStudy (anonymous):

Read the rules then. They say help and guide them through the answer it doesnt say if they ask for the answer just tell them it and only help if they ask to be guided through